Calculate the concentration of urea (NH2CONH2) that has an osmotic pressure of 7600torr at 25C.


1
Expert's answer
2021-04-09T04:25:53-0400

1 torr = 0.133 kPa

= 7600 × 0.133 = 1010.8Kpa

P= 5×0.0821×273×60/100

= 111.2×1000/1010.8

= 110.01187